Hy, this is from a book.
An investor has 3 shares, Delta Airlines, Hilton Hotels and Mcdonalds. He says that 2 days ago the shares went down 350$ and yesterday they went up 600$. His stock broker remembers that 2 days ago Delta lost 1$ per share, Hilton lost 1.5$, and Mc gained 0.5$. he also remembers that yesterday Delta gained 1.5$, Hilton lost again 0.5$ and Mc gained 1$. Explain why there isnt enough information to calculate the number of shares the investor has. But if the investor says that he has 200 shares of Mc, how many does he posses of the other 2?
I dont know how to put this info into 2 equations with 3 unknows. thanks i hope you can help me.

Answer by pushpaharan(47)

let x,y,z are no of shares in Delta Airlines, Hilton Hotels and Mcdonalds respectively.
2 days ago shares went doem by 350
therefore -1*x + (-1.5)*y + 0.5*z = -350
yesterday shares went up by 600
therefore 1.5*x + (-0.5)*y + 1*z = 600
you have 2 equation but 3 unknown, therefore cannot be solved.
since z=200
therfore
-x-1.5y+0.5*200 = -350
-x-1.5y = -350 - 100 = -450
x+1.5y = 450 ----(1)
other one
1.5x-0.5y+200 = 600
1.5x-0.5y=400
let's multiply the equation by 3
4.5x-1.5y = 1200 -----(2)
(1)+(2) => 5.5x = 1800
x = 327
sub x= 327 in (1)
327+1.5y=450
1.5y= 123
y=82